# The canonical lamination calibrated by a cohomology class

### v3

@@ A $k$-current of finite mass is a continuous linear functional on the space
of bounded continuous $d-k$-forms. [^3] Be warned: this convention agrees with currents in algebraic geometry, where currents are viewed as generalizations of forms, but not geometric measure theory, where currents are viewed as generalizations of submanifolds. We denote the action of a current $T$ on a form $\varphi$ by $\int_{M}T\wedge\varphi$.
@@
If $T$ represents a $d-k$-dimensional submanifold $\Sigma$, in the sense that $\int_{M}T\wedge F=\int_{\Sigma}F$, then $\mathbf{M}(T)=\mathrm{vol}(\Sigma)$. A function $u\in L^{1}_{\mathrm{loc}}(M)$ has bounded variation, denoted $u\in BV(M)$, if $\mathop{}\!\mathrm{d}u$ is a $1$-current of finite mass, in which case $\mathbf{M}(\mathop{}\!\mathrm{d}u)$ is the total variation of $u$, and we write $\int_{M}\star|\mathop{}\!\mathrm{d}u|$ to mean $\mathbf{M}(\mathop{}\!\mathrm{d}u)$.
@@
The next theorem is essentially [Anzellotti1983, Theorem 1.5], but we sketch the argument because Anzellotti did not formulate it in such generality.
@@
By the $BV$ Sobolev embedding theorem, [evans2015measure, §5.6], for every $\chi\in C^{\infty}_{\mathrm{cpt}}(M)$, $\chi u$ belongs to the dual space of $L^{d}(M,\Omega^{d-k})$. Therefore for every $\chi\in C^{\infty}_{\mathrm{cpt}}(M)$, $\langle\mathop{}\!\mathrm{d}u\wedge F,\chi\rangle$ is finite, so $\mathop{}\!\mathrm{d}u\wedge F$ is well-defined as a distribution.
@@
But, by the portmanteau theorem [kechris2012classical, Theorem 17.20],
@@ 2.3. Calibrated geometry
We recall calibrated geometry, which was developed by Harvey and Lawson [Harvey82].
@@
If $\Sigma$ is $F$-calibrated, then for any $k-1$-dimensional submanifold $\Lambda$,
@@
The definition of $F$-calibration extends to currents. If $F$ is a calibration $k$-form, a $d-k$-current $T$ is $F$-calibrated if
@@
By Anzellotti’s theorem, Theorem 2.4, this definition makes sense as long as $T$ has locally finite mass. If $T$ is $F$-calibrated, then for any $d-k-1$-current $S$, $\mathbf{M}(T)\leq\mathbf{M}(T+\mathop{}\!\mathrm{d}S)$.
@@
where $\operatorname{PD}(\omega)$ is the Poincaré dual of the class $\omega$, and $T$ ranges over $d-k$-currents. The costable norm $\|\cdot\|_{\infty}$ is the dual norm of $\|\cdot\|_{1}$ on $H^{k}(M,\mathbf{R})$. The following theorem is a special case of the main theorem of [Federer1974, §4] but it is essential to us, so we sketch the proof.
@@
For each representative $F$ of $\rho$, and with $T$ ranging over all $d-k$-currents,
@@
This implies that $\mathop{}\!\mathrm{d}F=0$ and $[F]=\rho$. ∎
@@ 2.4. Laminations
We use roughly the same formalism for laminations as in [Morgan88], which we also used in [BackusCML]. Let $M$ be a Riemannian manifold. Fix an interval $I\subset\mathbf{R}$ and a box $J\subset\mathbf{R}^{d-1}$. A (codimension-$1$, Lipschitz) laminar flow box is a Lipschitz coordinate chart $\Psi:I\times J\to M$ and a compact set $K\subseteq I$, called the local leaf space, such that for each $k\in K$, $\Psi|_{\{k\}\times J}$ is a $C^{1}$ embedding, and the leaf $\Psi(\{k\}\times J)$ is a $C^{1}$ complete hypersurface in $\Psi(I\times J)$. Two laminar flow boxes belong to the same laminar atlas if the transition map preserves the local leaf spaces.
@@
Theorem 2.8 ([BackusCML, Theorem A]).
Let $\mathcal{S}$ be a set of disjoint minimal hypersurfaces in $M$, such that $\bigcup_{N\in\mathcal{S}}N$ is a closed set, and $\sup_{N\in\mathcal{S}}\|\mathrm{I\!I}_{N}\|_{C^{0}}<\infty$. Then $\mathcal{S}$ is the set of leaves of a Lipschitz lamination $\lambda$, such that the normal vector to the leaves of $\lambda$ extends to a Lipschitz section of a line bundle on $M$.
In our application, $d\leq 7$ and the hypersurfaces in $\mathcal{S}$ are stable, so we can check the hypothesis on curvature in Theorem 2.8 using [Schoen81]. In my experience, it is a common misconception that the hypothesis on curvature can be removed, but the next example shows that it cannot be.
@@
An arbitrary lamination cannot be viewed as a current, but following Ruelle and Sullivan [Ruelle75], we view laminations which have been equipped with transverse measures and orientations as currents, so our next task is to define the Ruelle-Sullivan current.
@@
Let $(\lambda,\mu)$ be a measured oriented lamination. It is a straightforward modification of the arguments of [daskalopoulos2020transverse, §8] to show that the Ruelle-Sullivan current $T_{\mu}$ is a closed $1$-current which is well-defined, in the sense that $T_{\mu}$ does not depend on the choice of laminar atlas. Furthermore, by [BackusCML, Lemma 3.1],
@@
Let $\lambda$ be a lamination. A Borel set $E\subseteq\operatorname{supp}\lambda$ is saturated if, for every leaf $N$ of $\lambda$ such that $N\cap E$ is nonempty, $N\subseteq E$. Every leaf of $\lambda$ is Borel, and therefore saturated. A sublamination of $\lambda$ is a closed saturated set. Every sublamination of $\lambda$ the flow boxes of $\lambda$ and therefore is itself a lamination.
@@
We shall also need a form of the Morgan–Shelan decomposition, [Morgan88, Theorem I.3.2], of a measured lamination. To formulate it, let us say that a lamination $\lambda$ is exceptional [^4] Exceptional laminations are often called minimal, but that clashes with the use of the word “minimal” to refer to vanishing mean curvature, so we have not adopted this terminology. if every leaf of $\lambda$ is dense in $\operatorname{supp}\lambda$, and $\lambda$ is not a single closed leaf. A lamination $\lambda$ is a parallel family of closed leaves if there exists a closed leaf $N$ of $\lambda$ with trivial normal bundle, such that every leaf of $\lambda$ is a section of the normal bundle of $N$.
Theorem 2.12 (Morgan–Shelan decomposition).
@@
First observe that the proof of [Morgan88, Theorem I.3.2] goes through for any lamination $\lambda$ such that no leaf of $\lambda$ is dense in $M$, even if $\lambda$ is a foliation. It then remains to rule out the case that $\kappa$ is a family of sections of a nontrivial normal bundle of a closed leaf: this holds because $\lambda$ is oriented. ∎
@@ 3.1. Calibrated laminations
Since a current can be approximated by smooth $1$-forms in the weakstar topology on currents, every current has a cohomology class $[T]\in H^{1}(M,\mathbf{R})$. Thus, the homology class $[\lambda]\in H_{d-1}(M,\mathbf{R})$ is the Poincaré dual of $[T_{\lambda}]$.
@@
Let $\lambda$ be a measured oriented lamination, and assume that $M$ is compact. Then $\lambda$ is homologically minimizing, if for every measured oriented lamination $\kappa$ such that $[\lambda]=[\kappa]$,
@@
To upgrade $\lambda$ from an almost calibrated lamination to a calibrated lamination, we first, given $\sigma_{\alpha,k}$, choose $k_{j}$ such that $\sigma_{\alpha,k_{j}}$ is calibrated and $k_{j}\to k$. By Theorem 2.1, we can find a continuous $d-2$-form $A$ defined near $\sigma_{\alpha,k}$ with $F=\mathop{}\!\mathrm{d}A$. This justifies the following application of Stokes’ theorem:
@@ 3.2. Functions of least gradient
In either case we write $[\mathop{}\!\mathrm{d}u]=\alpha$, and write $\int_{M}\star|\mathop{}\!\mathrm{d}u|$ or $\mathbf{M}(\mathop{}\!\mathrm{d}u)$ to refer to the mass of the $1$-current that $\mathop{}\!\mathrm{d}u$ induces on $M$. If $[\mathop{}\!\mathrm{d}u]=0$ then we identify $u$ with the function that it induces on $M$.
@@
This sequence is bounded in $BV_{\mathrm{loc}}(\tilde{M},\mathbf{R})$, so by Alaoglu’s theorem, it has a subsequence which converges in the weakstar topology of $BV_{\mathrm{loc}}$ to some function $u$ such that $\mathbf{M}(\mathop{}\!\mathrm{d}u)\leq\|\alpha\|_{1}$. By testing $\mathop{}\!\mathrm{d}u_{n}$ against smooth $d-1$-forms on $M$, we see that $[\mathop{}\!\mathrm{d}u]=\alpha$ and so $u$ has least gradient. ∎
@@
If $u$ has least gradient, then [BackusCML, Theorem B] implies that there is a measured oriented lamination $\tilde{\lambda}_{u}$ of minimal hypersurfaces on $\tilde{M}$ whose leaves are level sets of $u$, and whose Ruelle-Sullivan current is $\mathop{}\!\mathrm{d}u$. Since $u$ is equivariant, $\tilde{\lambda}_{u}$ descends to a lamination $\lambda_{u}$ on $M$ such that $\mathbf{M}(\lambda_{u})=\mathbf{M}(\mathop{}\!\mathrm{d}u)$. Since $u$ has least gradient, $\lambda_{u}$ is homologically minimizing.
Conversely, if such a lamination exists, [BackusCML, Theorem B] implies that $u$ locally has least gradient and $\mathbf{M}(\mathop{}\!\mathrm{d}u)=\mathbf{M}(\lambda_{u})$, so $u$ has least gradient. ∎
@@ 3.3. Duality of calibrations and laminations
It is natural to ask if there is a continuous calibration in $\rho$, as was assumed in [bangert_cui_2017, Freedman_2016]. If $d=2$ one might try to generalize the argument of [Evans08] to obtain a Hölder continuous calibration, but if $d\geq 8$ then continuous calibrations need not exist [liu2023homologically]. The situation that $3\leq d\leq 7$ remains unclear. If $\mathrm{Ric}_{M}\geq 0$, then the Bochner argument shows that the harmonic representative of $\rho$ is a calibration; however, the Bochner argument actually shows that $M=\mathbf{S}^{1}\times N$ where $N$ is the calibrated hypersurface, so this is not very interesting.
In the setting of the Dirichlet problem for a domain on euclidean space, Mazón, Rossi, and Segura de León [Mazon14] proved that a $BV$ function has least gradient iff it is calibrated by some calibration. In fact, the same duality holds here, but in the equivariant setting the proof is trivial.
@@
Conversely, if $u$ has least gradient, then let $\alpha:=[\mathop{}\!\mathrm{d}u]$ and choose $\rho\in H^{d-1}(M,\mathbf{R})$ such that $\langle\rho,\alpha\rangle=\|\alpha\|_{1}$ and $\|\rho\|_{\infty}=1$. In particular, there exists a calibration $F$ such that $[F]=\rho$, and
@@
The next lemma was observed by Bangert and Cui, [bangert_cui_2017], in the setting that $F$ is continuous and we require no regularity on the laminations involved.
@@
defined on the open disk $U$ bounded by the circle $(x-2)^{2}+y^{2}=1$. Then $v$ is $\infty$-harmonic, meaning that
@@
To see this, it is best to work in polar coordinates, $x+iy=re^{i\theta}$. Then $v(re^{i\theta})=\theta$, so $\mathop{}\!\mathrm{d}v=\mathop{}\!\mathrm{d}\theta$. The euclidean metric is
@@
Also, $|\mathop{}\!\mathrm{d}\theta|=r^{-1}$, which only attains its maximum at the boundary point $x+iy=1$. In particular, $\|\mathop{}\!\mathrm{d}v\|_{L^{\infty}}=1$ and $\mathop{}\!\mathrm{d}v$ is a calibration on $U$. Since $v$ is $\infty$-harmonic, $v$ minimizes its Lipschitz constant, $\|\mathop{}\!\mathrm{d}v\|_{L^{\infty}}$, among all functions with the same boundary data [Crandall2008]. But if $u$ is a function on $U$ such that $\mathop{}\!\mathrm{d}u$ is $\mathop{}\!\mathrm{d}v$-calibrated, then
$\operatorname{supp}\mathop{}\!\mathrm{d}u\subseteq\{|\mathop{}\!\mathrm{d}v|=1\}\subset\partial U$
so $u$ is constant away from the boundary, hence is constant.
A more geometric way to visualize this phenomenon is to notice that the streamlines of $v$ – that is, the integral curves of the gradient of $v$ – are the circles centered on 00. If $u$ was $\mathop{}\!\mathrm{d}v$-calibrated, then the level sets of $u$ would correspond to the streamlines of $v$. However, since any $\mathop{}\!\mathrm{d}v$-calibrated function $u$ has least gradient, the level sets of $u$ must be straight lines.

### v4

@@ A $k$-current of finite mass is a continuous linear functional on the space
of bounded continuous dkd-k-forms. [^3] Be warned: this convention agrees with currents in algebraic geometry, where currents are viewed as generalizations of forms, but not geometric measure theory, where currents are viewed as generalizations of submanifolds. We denote the action of a current $T$ on a form $\varphi$ by $\int_{M}T\wedge\varphi$.
@@
If $T$ represents a dkd-k-dimensional submanifold $\Sigma$, in the sense that $\int_{M}T\wedge F=\int_{\Sigma}F$, then $\mathbf{M}(T)=\mathrm{vol}(\Sigma)$. A function $u\in L^{1}_{\mathrm{loc}}(M)$ has bounded variation, denoted $u\in BV(M)$, if $\mathop{}\!\mathrm{d}u$ is a $1$-current of finite mass, in which case $\mathbf{M}(\mathop{}\!\mathrm{d}u)$ is the total variation of $u$, and we write $\int_{M}\star|\mathop{}\!\mathrm{d}u|$ to mean $\mathbf{M}(\mathop{}\!\mathrm{d}u)$.
@@
The next theorem is essentially [1, Theorem 1.5], but we sketch the argument because Anzellotti did not formulate it in such generality.
@@
By the $BV$ Sobolev embedding theorem, [21, §5.6], for every $\chi\in C^{\infty}_{\mathrm{cpt}}(M)$, $\chi u$ belongs to the dual space of $L^{d}(M,\Omega^{d-k})$. Therefore for every $\chi\in C^{\infty}_{\mathrm{cpt}}(M)$, $\langle\mathop{}\!\mathrm{d}u\wedge F,\chi\rangle$ is finite, so $\mathop{}\!\mathrm{d}u\wedge F$ is well-defined as a distribution.
@@
But, by the portmanteau theorem [32, Theorem 17.20],
@@ 2.3. Calibrated geometry
We recall calibrated geometry, which was developed by Harvey and Lawson [30].
@@
If $\Sigma$ is $F$-calibrated, then for any k1k-1-dimensional submanifold $\Lambda$,
@@
The definition of $F$-calibration extends to currents. If $F$ is a calibration $k$-form, a dkd-k-current $T$ is $F$-calibrated if
@@
By Anzellotti’s theorem, Theorem 2.4, this definition makes sense as long as $T$ has locally finite mass. If $T$ is $F$-calibrated, then for any dk1d-k-1-current $S$, $\mathbf{M}(T)\leq\mathbf{M}(T+\mathop{}\!\mathrm{d}S)$.
@@
where $\operatorname{PD}(\omega)$ is the Poincaré dual of the class $\omega$, and $T$ ranges over dkd-k-currents. The costable norm $\|\cdot\|_{\infty}$ is the dual norm of $\|\cdot\|_{1}$ on $H^{k}(M,\mathbf{R})$. The following theorem is a special case of the main theorem of [22, §4] but it is essential to us, so we sketch the proof.
@@
For each representative $F$ of $\rho$, and with $T$ ranging over all dkd-k-currents,
@@
This implies that $\mathop{}\!\mathrm{d}F=0$ and $F]=\rho$. ∎
@@ 2.4. Laminations
We use roughly the same formalism for laminations as in [40], which we also used in [6]. Let $M$ be a Riemannian manifold. Fix an interval $I\subset\mathbf{R}$ and a box $J\subset\mathbf{R}^{d-1}$. A (codimension-$1$, Lipschitz) laminar flow box is a Lipschitz coordinate chart $\Psi:I\times J\to M$ and a compact set $K\subseteq I$, called the local leaf space, such that for each $k\in K$, $\Psi|_{\{k\}\times J}$ is a $C^{1}$ embedding, and the leaf $\Psi(\{k\}\times J)$ is a $C^{1}$ complete hypersurface in $\Psi(I\times J)$. Two laminar flow boxes belong to the same laminar atlas if the transition map preserves the local leaf spaces.
@@
Theorem 2.8 ([6, Theorem A]).
Let $\mathcal{S}$ be a set of disjoint minimal hypersurfaces in $M$, such that $\bigcup_{N\in\mathcal{S}}N$ is a closed set, and $\sup_{N\in\mathcal{S}}\|\mathrm{I\!I}_{N}\|_{C^{0}}<\infty$. Then $\mathcal{S}$ is the set of leaves of a Lipschitz lamination $\lambda$, and the normal vector to the leaves of $\lambda$ extends to a Lipschitz section of a line bundle over $M$.
In our application, $d\leq 7$ and the hypersurfaces in $\mathcal{S}$ are stable, so we can check the hypothesis on curvature in Theorem 2.8 using [44]. In my experience, it is a common misconception that the hypothesis on curvature can be removed, but the next example shows that it cannot be.
@@
An arbitrary lamination cannot be viewed as a current, but following Ruelle and Sullivan [42], we view laminations which have been equipped with transverse measures and orientations as currents, so our next task is to define the Ruelle-Sullivan current.
@@
Let $(\lambda,\mu)$ be a measured oriented lamination. It is a straightforward modification of the arguments of [17, §8] to show that the Ruelle-Sullivan current $T_{\mu}$ is a closed $1$-current which is well-defined, in the sense that $T_{\mu}$ does not depend on the choice of laminar atlas. Furthermore, by [6, Lemma 3.1],
@@
Let $\lambda$ be a lamination. A Borel set $E\subseteq\operatorname{supp}\lambda$ is saturated if, for every leaf $N$ of $\lambda$ such that $N\cap E$ is nonempty, $N\subseteq E$. Every leaf of $\lambda$ is Borel, and therefore saturated. A sublamination of $\lambda$ is a closed saturated set. Every sublamination of $\lambda$ inherits the flow boxes of $\lambda$ and therefore is itself a lamination.
@@
We shall also need a form of the Morgan–Shalen decomposition, [40, Theorem I.3.2], of a measured lamination. To formulate it, let us say that a lamination $\lambda$ is exceptional [^4] Exceptional laminations are often called minimal, but that clashes with the use of the word “minimal” to refer to vanishing mean curvature, so we have not adopted this terminology. if every leaf of $\lambda$ is dense in $\operatorname{supp}\lambda$, and $\lambda$ is not a single closed leaf. A lamination $\lambda$ is a parallel family of closed leaves if there exists a closed leaf $N$ of $\lambda$ with trivial normal bundle, such that every leaf of $\lambda$ is a section of the normal bundle of $N$.
Theorem 2.12 (Morgan–Shalen decomposition).
@@
First observe that the proof of [40, Theorem I.3.2] goes through for any lamination $\lambda$ such that no leaf of $\lambda$ is dense in $M$, even if $\lambda$ is a foliation. It then remains to rule out the case that $\kappa$ is a family of sections of a nontrivial normal bundle of a closed leaf: this holds because $\lambda$ is oriented. ∎
2.5. Perimeter-minimizing sets
In this section only, $M$ denotes a complete Riemannian manifold of bounded curvature; we do not assume that $d=7$ or $M$ is closed. A Borel set $U\subseteq M$ is perimeter-minimizing if for every ball $B\subseteq M$ and every Borel set $V\subseteq M$ such that $1_{U}=1_{V}$ on $\partial B$, $\mathrm{vol}(\partial U)\leq\mathrm{vol}(\partial V)$ in the sense of currents. In the proof of Theorem 1.6(2), we shall need an estimate on perimeter-minimizing sets, which we now prove; the reader can ignore this section until then. See [25, Chapter 5] for the proof when $M$ is an open subset of euclidean space.
Lemma 2.13.
$\mathrm{vol}(U\cap B(x,r))\geq cr^{d}.$ (2.4)
Proof.
where $c_{d}>0$ is the euclidean isoperimetric constant. We can reason as in the proof of [25, Proposition 5.14] to see that for almost every $0<\rho<r$,
Let $f(\rho):=\mathrm{vol}(U\cap B(x,\rho))^{1/d}$, so that
Therefore $f(\rho)\geq\rho/(4dc_{d})$, as desired. ∎
Lemma 2.14.
Proof.
@@ 3.1. Calibrated laminations
Since a current can be approximated by smooth $1$-forms in the weakstar topology on currents, every current has a cohomology class $T]\in H^{1}(M,\mathbf{R})$. Thus, the homology class $\lambda]\in H_{d-1}(M,\mathbf{R})$ is the Poincaré dual of $T_{\lambda}]$.
@@
Let $\lambda$ be a measured oriented lamination, and assume that $M$ is compact. Then $\lambda$ is homologically minimizing, if for every measured oriented lamination $\kappa$ such that $\lambda]=[\kappa]$,
@@
To upgrade $\lambda$ from an almost calibrated lamination to a calibrated lamination, we first, given $\sigma_{\alpha,k}$, choose $k_{j}$ such that $\sigma_{\alpha,k_{j}}$ is calibrated and $k_{j}\to k$. By Theorem 2.1, we can find a continuous d2d-2-form AA defined near $\sigma_{\alpha,k}$ with $F=\mathop{}\!\mathrm{d}A$. This justifies the following application of Stokes’ theorem:
@@ 3.2. Functions of least gradient
In either case we write $\mathop{}\!\mathrm{d}u]=\alpha$, and write $\int_{M}\star|\mathop{}\!\mathrm{d}u|$ or $\mathbf{M}(\mathop{}\!\mathrm{d}u)$ to refer to the mass of the $1$-current that $\mathop{}\!\mathrm{d}u$ induces on $M$. If $\mathop{}\!\mathrm{d}u]=0$ then we identify $u$ with the function that it induces on $M$.
@@
This sequence is bounded in $BV_{\mathrm{loc}}(\tilde{M},\mathbf{R})$, so by Alaoglu’s theorem, it has a subsequence which converges in the weakstar topology of $BV_{\mathrm{loc}}$ to some function $u$ such that $\mathbf{M}(\mathop{}\!\mathrm{d}u)\leq\|\alpha\|_{1}$. By testing $\mathop{}\!\mathrm{d}u_{n}$ against smooth d1d-1-forms on $M$, we see that $\mathop{}\!\mathrm{d}u]=\alpha$ and so $u$ has least gradient. ∎
@@
If $u$ has least gradient, then [6, Theorem B] implies that there is a measured oriented lamination $\tilde{\lambda}_{u}$ of minimal hypersurfaces on $\tilde{M}$ whose leaves are level sets of $u$, and whose Ruelle-Sullivan current is $\mathop{}\!\mathrm{d}u$. Since $u$ is equivariant, $\tilde{\lambda}_{u}$ descends to a lamination $\lambda_{u}$ on $M$ such that $\mathbf{M}(\lambda_{u})=\mathbf{M}(\mathop{}\!\mathrm{d}u)$. Since $u$ has least gradient, $\lambda_{u}$ is homologically minimizing.
Conversely, if such a lamination exists, [6, Theorem B] implies that $u$ locally has least gradient and $\mathbf{M}(\mathop{}\!\mathrm{d}u)=\mathbf{M}(\lambda_{u})$, so $u$ has least gradient. ∎
@@ 3.3. Duality of calibrations and laminations
It is natural to ask if there is a continuous calibration in $\rho$, as was assumed in [8, 23]. If $d=2$ one might try to generalize the argument of [20] to obtain a Hölder continuous calibration, but if $d\geq 8$ then continuous calibrations need not exist [35]. The situation that $3\leq d\leq 7$ remains unclear. If $\mathrm{Ric}_{M}\geq 0$, then the Bochner argument shows that the harmonic representative of $\rho$ is a calibration; however, the Bochner argument actually shows that $M=\mathbf{S}^{1}\times N$ where $N$ is the calibrated hypersurface, so this is not very interesting.
In the setting of the Dirichlet problem for a domain on euclidean space, Mazón, Rossi, and Segura de León [39] proved that a $BV$ function has least gradient iff it is calibrated by some calibration. In fact, the same duality holds here, but in the equivariant setting the proof is trivial.
@@
Conversely, if $u$ has least gradient, then let $\alpha:=[\mathop{}\!\mathrm{d}u]$ and choose $\rho\in H^{d-1}(M,\mathbf{R})$ such that $\langle\rho,\alpha\rangle=\|\alpha\|_{1}$ and $\|\rho\|_{\infty}=1$. In particular, there exists a calibration $F$ such that $F]=\rho$, and
@@
The next lemma was observed by Bangert and Cui, [8], in the setting that $F$ is continuous and we require no regularity on the laminations involved.
@@
defined on the open disk $U\subset\mathbf{C}$ bounded by the circle $(x-2)^{2}+y^{2}=1$. Working in the polar coordinates $x+iy=re^{i\theta}$, one has $\mathop{}\!\mathrm{d}v=\mathop{}\!\mathrm{d}\theta$, so $|\mathop{}\!\mathrm{d}v|=r^{-1}$, which attains its maximum only at the boundary point $x+iy=1$, where it is $1$. Therefore $\mathop{}\!\mathrm{d}v$ is a calibration.
Let $\iota:\partial U\to\overline{U}$ be the inclusion map. If $F$ is another calibration which satisfies the boundary condition $\iota^{*}F=\iota^{*}\mathop{}\!\mathrm{d}v$, then since $\mathop{}\!\mathrm{d}F=0$, we can write $F=\mathop{}\!\mathrm{d}w$. Then the trace of wvw-v along $\partial U$ is a constant, which we may assume to be 0. But $v$ is $\infty$-harmonic, meaning that
@@
Indeed, in polar coordinates, the euclidean metric becomes
@@
Since $v$ is $\infty$-harmonic, $\|F\|_{L^{\infty}}\geq\|\mathop{}\!\mathrm{d}v\|_{L^{\infty}}$ [15], so that the “costable norm” of the boundary data $\iota^{*}\mathop{}\!\mathrm{d}v$ is $1$.
But if $u$ is a function on $U$ such that $\mathop{}\!\mathrm{d}u$ is $\mathop{}\!\mathrm{d}v$-calibrated, then
$\operatorname{supp}\mathop{}\!\mathrm{d}u\cap U\subseteq\{|\mathop{}\!\mathrm{d}v|=1\}\cap U=\emptyset,$
so $u$ is constant! A more geometric way to visualize this phenomenon is to notice that the streamlines of $v$ – that is, the integral curves of the gradient of $v$ – are the circles centered on 0. If $u$ was $\mathop{}\!\mathrm{d}v$-calibrated and $\gamma$ was a level curve of $u$, then $\gamma$ would be a streamline of $v$, but would also be $\mathop{}\!\mathrm{d}v$-calibrated. We would conclude that circles are geodesics, which is a contradiction.